I am having a difficult time solving this equation:
3(x+8) = 7x -2

Answer by longjonsilver(2297)

3(x+8) = 7x -2
3x + 24 = 7x - 2
24 = 4x - 2
26 = 4x or... 4x = 26

x = 26/4
x = 6.5

jon.